Consider the equation %ΔM + %ΔV ≈ %ΔP + %ΔY. If there is no change in real GDP (%ΔY = 0) and there is no change in the money supply (%ΔM = 0), what would have to happen for there to be inflation (%ΔP > 0)? Money would have to change hands less often, %ΔV < 0. Money supply times money velocity would have to go down, %ΔM + %ΔV < 0. Money would have to change hands more often, %ΔV > 0. Nominal GPD would have to go down, %ΔP + %ΔY < 0. The money velocity would have to remain the same, %ΔV = 0.
Answer #1

Solution :

We know that,

M + V = P + Y

So, from the above question, it is said that there is no change in real GDP and money supply.

So, from M + V = P + Y , this equation we know that any changes in money supply would be equivalent to change in inflation rate. Hence, the correct answer would be,

The money velocity would have to remain the same,

% change in V = 0.